# Long-Context Language Modeling with Parallel Context Encoding

### v1

@@ Abstract
Extending large language models (LLMs) to process longer inputs is crucial for numerous applications. However, the considerable computational cost of transformers, coupled with limited generalization of positional encoding, restricts the size of their context window. We introduce Context Expansion with Parallel Encoding (CEPE ), a framework that can be applied to any existing decoder-only LLMs to extend their context window. CEPE adopts a small encoder to process long inputs chunk by chunk and enables the frozen decoder to leverage additional contexts via cross-attention. CEPE is efficient, generalizable, and versatile: trained with 8K-token documents, CEPE extends the context window of LLaMA-2 to 128K tokens, offering 10×10$\times$10 × the throughput with only 1/6 of the memory. CEPE yields strong performance on language modeling and in-context learning. CEPE also excels in retrieval-augmented applications, while existing long-context models degenerate with retrieved contexts. We further introduce a CEPE variant that can extend the context window of instruction-tuned models with only unlabeled data, and showcase its effectiveness on LLaMA-2-Chat, leading to a strong instruction-following model that can leverage very long context on downstream tasks. [^1] https://github.com/princeton-nlp/CEPE.
@@
Figure 1: A comparison between CEPE and other techniques of extending LLMs’ context window, including YaRN Peng et al. (2023), StreamingLLM Xiao et al. (2023), and RePlug Shi et al. (2023b). CEPE trained on 8K tokens can generalize to 128K tokens with minimal computational and memory costs.
@@ 1 Introduction
Long and extensible context is crucial for large language models (LLMs) to effectively perform complex tasks, such as summarizing a book or answering questions with hundreds of retrieved web pages. However, several challenges limit the ability of LLMs to leverage long context: (1) LLMs and popular positional encodings Raffel et al. (2020); Su et al. (2021) do not generalize to sequence lengths longer than the lengths seen during training Press et al. (2022), even after additional fine-tuning (Together, 2023a; Chen et al., 2023a, b; Peng et al., 2023, inter alia). (2) Transformers Vaswani et al. (2017)—the predominant architecture of LLMs—incur a quadratic computational cost and a linear memory cost with respect to the input length, making it prohibitive to use for long sequences. (3) High-quality long-context data, such as long instruction-following data, are scarce and difficult to obtain Wang et al. (2023); Xiong et al. (2023).
@@
Besides directly fine-tuning LLMs on longer inputs, a series of inference-time modification methods have been proposed recently to scale up the effective context window, either by modifying the attention mechanism Bertsch et al. (2023); Xiao et al. (2023); Ivgi et al. (2023) or encoding chunks of context in separate forward passes Shi et al. (2023b); Ratner et al. (2023); Lin et al. (2023). While these methods generalize to longer sequences, the model often fails to effectively leverage the extra tokens and can incur greater inference costs.
@@
Figure 2: The CEPE architecture. The encoder model encodes the additional 3 chunks ($k=3$) of context $\mathcal{C}$ in parallel, and the final hidden representations from the encoder model are concatenated and used as inputs to the cross-attention layers in the decoder model. The cross-attention layers attend to the encoder representations between the self-attention and feed-forward layers in the decoder model.
@@
In this work, we propose an efficient and lightweight solution to extending the context window of LLMs, called Context Expansion with Parallel Encoding (CEPE). CEPE is applicable to any pre-trained decoder-only LM by adding two components: (1) a small encoder that encodes the long context in chunks, and (2) a cross-attention module that is inserted at each layer of the decoder to attend to the encoder representations (Figure 2). With a careful selection of unlabeled training data, CEPE can leverage not only long-context documents, but also retrieved documents effectively.
@@
CEPE offers several benefits: (1) Length generalization: CEPE is not limited by positional encoding constraints as the long context is encoded in segments, each with its own positional encoding. (2) Efficiency: using a small encoder and processing context in parallel reduce computational cost. Since cross-attention attends only to the last layer’s representations from the encoder, CEPE requires much less memory compared to decoder-only LMs, which cache the key-value pairs of every token in every layer. (3) Reduced training cost: unlike full fine-tuning approaches, we only tune the encoder and the cross-attention while keeping the large decoder LM frozen; augmenting a 7B decoder with a 400M encoder and cross-attention layers (1.4B parameters) can be done with one A100 80GB GPU.
@@
We apply CEPE to LLaMA-2 and train it on a filtered version of RedPajama (Together, 2023b) for 20B tokens—only $1$% of the pre-training budget of LLaMA-2. We first show that CEPE-LLaMA-2, trained with 8K input length, continues to improve perplexity on longer input up to 128K tokens. Then, we apply CEPE to a retrieval-augmented setting, as our larger context window allows incorporating more retrieved documents. Compared to existing methods, CEPE achieves better performance on both retrieval-augmented language modeling and open-domain question answering. Additionally, we demonstrate that CEPE can effectively leverage more demonstrations for in-context learning Brown et al. (2020). All the above is achieved with a much lower memory and computational cost than most previous solutions.
@@
Finally, we propose CEPE-Distilled (CEPED), which extends the context window of instruction-tuned models, using only unlabeled data. CEPED distills the behavior of the original instruction-tuned model to the new architecture through an auxiliary KL divergence loss, which eliminates the need to curate expensive long-context instruction-following data Ivison et al. (2023). We apply CEPED to LLaMA-2-Chat Touvron et al. (2023b) and show that while preserving their instruction understanding ability, CEPED-LLaMA-2-Chat can incorporate more context and improve performance on long-text understanding tasks  (Shaham et al., 2023).
@@
To conclude, CEPE is a lightweight framework that can extend context windows of any pre-trained or instruction-tuned LMs. We hope CEPE can empower future LLM research with cheap and effective long-context abilities.
2 Our Method: CEPE
@@
We design CEPE to adapt pre-trained LLMs to perform long-context language modeling on sequences with many tokens (e.g., books). For retrieval augmentation, these long contexts may contain a set of retrieved passages instead. We first describe how CEPE modifies the architecture of LLMs to attend to representations encoded by a small encoder, and then describe how the CEPE modules are trained. Finally, we extend CEPE to CEPED, which extends the context window of instruction-tuned models using only unlabeled data.
@@ 2.1 Architecture
Given an input context with $T$ tokens $x_{1},...,x_{T}$, we consider the first $m$ tokens $x_{1},...,x_{m}$ as the additional context $\mathcal{C}$ and the last $n=T-m$ tokens $x_{m+1},...,x_{T}$ as the main input $X$. The additional context is split into chunks $\mathcal{C}=C_{1},...,C_{k}$, which can contain either segments of long documents or a set of retrieved passages. We use $\mathcal{M}_{\textrm{enc}}$ to denote the encoder model with hidden dimension $d_{\textrm{enc}}$ and $\mathcal{M}_{\textrm{dec}}$ to denote the decoder-only LLM with hidden dimension $d_{\textrm{dec}}$.
@@
where $\phi_{i}\in\mathbb{R}^{|C_{i}|\times d_{\textrm{dec}}}$ is the token-wise last layer hidden state from $\mathcal{M}_{\textrm{enc}}$ and $\Phi\in\mathbb{R}^{m\times d_{\textrm{dec}}}$. Note that $\mathcal{M}_{\textrm{enc}}$ is bidirectional; leveraging a bidirectional encoder results in more information-rich representations compared to unidirectional ones.
@@
In every decoder layer of the transformer  (Vaswani et al., 2017), we insert a cross-attention module between the self-attention and feed-forward layers. To construct the cross-attention module, we provide $\Phi$ as keys and values, and the hidden states of $X$ as queries. Note that in order for $\mathcal{M}_{\textrm{dec}}$ to attend to $\Phi$, the key and value projection matrices in the cross-attention module also serve as an up-projection that transforms the $d_{\textrm{enc}}$-dimensional $\Phi$ into a $d_{\textrm{dec}}$-dimensional embedding. Figure 2 illustrates the architecture of CEPE.
@@ 2.3 Training
We first pre-train a bidirectional masked-language model (MLM) on the RedPajama dataset. $\mathcal{M}_{\textrm{enc}}$ follows the configuration of RoBERTa-large Liu et al. (2019b) but shares the vocabulary with $\mathcal{M}_{\textrm{dec}}$, LLaMA-2. Future works applying CEPE to LLaMA-2 models can simply use our pre-trained encoder. We train $\mathcal{M}_{\textrm{enc}}$ for 100K steps with a batch size of $2,048$ and a sequence length of $512$ tokens. For more details, please refer to §A.1.
@@
We freeze the original weights of the decoder model and train only the added cross-attention layers as well as fine-tuning $\mathcal{M}_{\textrm{enc}}$ using the cross-entropy loss. We first adopt a warmup training stage designed to teach $\mathcal{M}_{\textrm{dec}}$ to copy from $\mathcal{M}_{\textrm{enc}}$ through cross-attention—for each position $i\leq T$, the objective is to generate $x_{i+1}$, conditioned on $\mathcal{M}_{\textrm{enc}}(x_{1},\ldots,x_{T})$ and $\mathcal{M}_{\textrm{dec}}(x_{1},\ldots,x_{i})$—i.e., $\mathcal{M}_{\textrm{enc}}$ and $\mathcal{M}_{\textrm{dec}}$ share the same input. The warmup stage uses 131M tokens sampled from the training set of RP. [^2] Preliminary experiments suggest this stage can stabilize training; ablations can be found in §7.1.
@@ 2.4 CEPED for Instruction-Tuned Models
Table 1: Long-context language modeling text perplexity on ArXiv and Book from RedPajama, PG19, ProofPile, and CodeParrot. Throughput compares the speed (number of tokens/second) of each model with that of LLaMA-2. All experiments are conducted on one A100 80GB GPU, except for LLaMA-2-32K and YaRN with 128K tokens, which require model parallelism, and are conducted on four A100 GPUs.
@@ 3 Long-context Language Modeling
We focus on 7B parameters model. Our baseline includes LLaMA-2 and its long-sequence fine-tuned versions: LLaMA-2-32K Together (2023a), YaRN-64K, and YaRN-128K Peng et al. (2023). LLaMA-2-32K was trained on RP with upsampled ArXiv and Books domains, and a mixture of other data , and YaRN was trained on PG19. While YaRN-64K and YaRN-128K were fine-tuned for only 1.7B and 3.4B tokens, respectively, fine-tuning all parameters requires much more memory relative to CEPE. We also evaluate on training-free long-context methods: StreamingLLM Xiao et al. (2023) and RePlug Shi et al. (2023a) with LLaMA-2-7B. Although RePlug was originally evaluated in a retrieval-augmented setting, we found that it also works well in long-context modeling, when viewing the long context as retrieved context. See §B for more implementation details.
@@ Results.
We show the results in Table 1. Compared to the two fully fine-tuned models, LLaMA-2-32K and YaRN-64K, CEPE achieves either lower or comparable perplexity across all datasets with lower memory usage and higher throughput. Furthermore, CEPE continues to improve on perplexity while maintaining low memory use at 128K tokens, well beyond its training lengths (8K); on the other hand, LLaMA-2-32K and YaRN-64K cannot generalize beyond its training length and the memory cost increases significantly. At 128K tokens, we outperform or achieve comparable perplexity with all applicable baselines on all domains.
@@ 4.2 Open-domain Question Answering
${}^{\dagger}$ 40 94.3 98.7 74.7 52.3 87.7 54.8 95.1 76.7 62.1 50.4 72.0
Table 3: ICL results averaged across 3 seeds. $k$ is the number of demonstrations. All models use 2 demonstrations in the decoder, and we add $+k^{\prime}$ demonstrations to the encoder for CEPE. ${}^{\dagger}$ denotes the oracle setting with $k=40$ demonstration in the decoder.
@@
Table 4: ZeroScrolls validation results. The total number of tokens includes both the input and generated tokens. For NarrativeQA (NQA) and Qspr (QASPER), we report the F1 scores. For QALT (QuALITY), we report accuracy. For GovReport (GvRp), SummScreenFD (SSFD), and QMSum, we report the ROUGE-L scores. CEPED uses 2K tokens in the decoder, and additional tokens are inputted through the encoder.
@@ Results.
The results are shown in Figure 3 CEPE consistently outperforms all models across all datasets and $k$. Notably, CEPE outperforms LLaMA-2-32K on NQ and TriviaQA by over 3 and 4 points in exact match, respectively. Furthermore, CEPE does not degrade in performance as the $k$ increases, while other models often perform worse at larger $k$, as they are sensitive to the large amount of redundant or irrelevant passages.
@@
Table 4 shows that CEPED improves upon LLaMA-2-Chat with 2K tokens across all tasks. The performance of CEPED improves or remains consistent as we scale up the number of tokens in the context window. Notably, CEPED improves upon LLaMA-2-Chat by 3 points in F1 scores on NQA, which has the longest average number of words among all ZeroScrolls tasks. Furthermore, CEPED outperforms LLaMA-2-32K Instruct on 4 out of the 6 tasks, despite being trained on unlabeled data. We provide qualitative examples and analysis in §C.3.

### v2

@@ Abstract
Extending large language models (LLMs) to process longer inputs is crucial for a wide range of applications. However, the substantial computational cost of transformers and limited generalization of positional encoding restrict the size of their context window. We introduce Context Expansion with Parallel Encoding (CEPE [Uncaptioned image]), a framework that can be applied to any existing decoder-only LLMs to extend their context window. CEPE employs a small encoder to process long inputs chunk by chunk, enabling the frozen decoder to utilize additional contexts via cross-attention. CEPE is efficient, generalizable, and versatile: trained with 8K-token documents, it extends the context window of LLAMA-2 to 128K tokens, offering 10× the throughput with only 1/6 of the memory. CEPE yields strong performance on language modeling and in-context learning. CEPE also excels in retrieval-augmented applications, while existing long-context models degenerate with retrieved contexts. We further introduce a CEPE variant that can extend the context window of instruction-tuned models using only unlabeled data, and showcase its effectiveness on LLaMA-2-Chat, leading to a strong instruction-following model that can leverage very long contexts on downstream tasks. [^1] Code and models are available at https://github.com/princeton-nlp/CEPE.
@@
Figure 1: A comparison between CEPE and other techniques of extending LLMs’ context window, including YaRN Peng et al. (2024), StreamingLLM Xiao et al. (2024b), and RePlug Shi et al. (2024). CEPE trained on 8K tokens can generalize to 128K tokens with minimal computational and memory costs.
@@ 1 Introduction
Enabling long and extensible context is crucial for large language models (LLMs) to effectively perform complex tasks, such as summarizing a book or answering questions with hundreds of retrieved Web pages. However, several challenges limit the ability of LLMs to leverage long context: (1) LLMs and popular positional encodings Raffel et al. (2020); Su et al. (2021) do not generalize to sequence lengths longer than the lengths seen during training Press et al. (2022), even after additional fine-tuning (Chen et al., 2023, 2024; Peng et al., 2024, inter alia). (2) Transformers Vaswani et al. (2017)—the predominant architecture of LLMs—incur a quadratic computational cost and a linear memory cost with respect to the input length, making it expensive to use for long sequences. (3) High-quality long-context data, such as long instruction-following data, are scarce and difficult to obtain Wang et al. (2023); Xiong et al. (2023).
@@
Besides directly fine-tuning LLMs on longer inputs, a series of inference-time modification methods have been proposed recently to scale up the effective context window, either by modifying the attention mechanism Bertsch et al. (2023); Xiao et al. (2024b); Ivgi et al. (2023) or encoding chunks of context in separate forward passes Shi et al. (2024); Ratner et al. (2023); Lin et al. (2024). While these methods generalize to longer sequences, the model often fails to effectively leverage the extra tokens and can incur greater inference costs.
@@
Figure 2: The CEPE Refer to caption architecture. The encoder model encodes the additional 3 chunks ($k=3$) of context $\mathcal{C}$ in parallel, and the final hidden representations from the encoder model are concatenated and used as inputs to the cross-attention layers in the decoder model. The cross-attention layers attend to the encoder representations between the self-attention and feed-forward layers in the decoder model.
@@
In this work, we propose an efficient and lightweight solution to extending the context window of LLMs, called Context Expansion with Parallel Encoding (CEPE [Uncaptioned image]). CEPE is applicable to any pre-trained decoder-only LM by adding two components: (1) a small encoder that encodes the long context in chunks, and (2) a cross-attention module that is inserted at each layer of the decoder to attend to the encoder representations (Figure 2). With a careful selection of unlabeled training data, CEPE can leverage not only long-context documents but also retrieved documents effectively.
@@
CEPE offers several benefits: (1) length generalization: CEPE is not limited by positional encoding constraints as the long context is encoded in chunks, each with its own positional encoding; (2) efficiency: using a small encoder and processing contexts in parallel reduce computational cost. Since cross-attention attends only to the last layer’s representations from the encoder, CEPE requires much less memory compared to decoder-only LMs, which cache the key-value pairs of every token in every layer; (3) reduced training cost: unlike full fine-tuning approaches, we only tune the encoder and the cross-attention while keeping the large decoder LM frozen; augmenting a 7B decoder with a 400M encoder and cross-attention layers (1.4B parameters) can be done with one A100 80GB GPU.
@@
We apply CEPE to LLaMA-2 Touvron et al. (2023b) and train it on a filtered version of RedPajama (Together, 2023b) for 20B tokens—only $1$% of the pre-training budget of LLaMA-2. We first show that CEPE-LLaMA-2, trained with 8K input length, continues to improve perplexity on longer input up to 128K tokens. Then, we apply CEPE to a retrieval-augmented setting, as our larger context window allows incorporating more retrieved documents. Compared to existing methods, CEPE achieves better performance on both retrieval-augmented language modeling and open-domain question answering. Additionally, we demonstrate that CEPE can effectively leverage more demonstrations for in-context learning Brown et al. (2020). All the above is achieved with a much lower memory and computational cost than most previous solutions.
@@
Finally, we propose CEPE-Distilled (CEPED), which extends the context window of instruction-tuned models, using only unlabeled data. CEPED distills the behavior of the original instruction-tuned model to the new architecture through an auxiliary KL divergence loss, which eliminates the need to curate expensive long-context instruction-following data Ivison et al. (2023). We apply CEPED to LLaMA-2-Chat and show that while preserving their instruction understanding ability, CEPED-LLaMA-2-Chat can incorporate more context and improve performance on long-text understanding tasks  (Shaham et al., 2023).
@@
To conclude, CEPE is a lightweight framework that can extend context windows of any base or instruction-tuned LMs. We hope CEPE can empower future LLM research with cheap and effective long-context abilities.
2 Method: CEPE [Uncaptioned image]
@@
We design CEPE to adapt pre-trained LLMs to perform long-context language modeling on sequences with many tokens (e.g., books). For retrieval augmentation, these long contexts may contain a set of retrieved passages instead. We first describe how CEPE modifies the architecture of LLMs to attend to representations encoded by a small encoder, and then describe how the CEPE modules are trained. Finally, we extend CEPE to CEPED, which expands the context window of instruction-tuned models using only unlabeled data.
@@ 2.1 Architecture
Given an input context with $T$ tokens $x_{1},...,x_{T}$, we consider the first $m$ tokens $x_{1},...,x_{m}$ as the additional context $\mathcal{C}$ and the last $n=T-m$ tokens $x_{m+1},...,x_{T}$ as the main input $X$. The additional context is split into chunks $\mathcal{C}=C_{1},...,C_{k}$, which can contain either segments within a long document or a set of retrieved passages. We use $\mathcal{M}_{\textrm{enc}}$ to denote the encoder model with hidden dimension $d_{\textrm{enc}}$ and $\mathcal{M}_{\textrm{dec}}$ to denote the decoder-only LLM with hidden dimension $d_{\textrm{dec}}$.
@@
where $\phi_{i}\in\mathbb{R}^{|C_{i}|\times d_{\textrm{dec}}}$ is the token-wise last layer hidden state from $\mathcal{M}_{\textrm{enc}}$ and $\Phi\in\mathbb{R}^{m\times d_{\textrm{dec}}}$. Note that $\mathcal{M}_{\textrm{enc}}$ is bidirectional, which results in more information-rich representations compared to unidirectional ones. While we do not preserve global positions across different chunks, experiments show that CEPE achieves better or comparable performance to full-attention models that do.
@@
In every decoder layer of the transformer, we insert a cross-attention module between the self-attention and feed-forward layers. To construct the cross-attention module, we provide $\Phi$ as keys and values, and the hidden states of $X$ as queries. Note that in order for $\mathcal{M}_{\textrm{dec}}$ to attend to $\Phi$, the key and value projection matrices in the cross-attention module also serve as an up-projection that transforms the $d_{\textrm{enc}}$-dimensional $\Phi$ into a $d_{\textrm{dec}}$-dimensional embedding. Figure 2 illustrates the architecture of CEPE.
@@ 2.3 Training
We first pre-train a bidirectional masked-language model on the RedPajama dataset. $\mathcal{M}_{\textrm{enc}}$ follows the configuration of RoBERTa-large Liu et al. (2019b) but shares the vocabulary with $\mathcal{M}_{\textrm{dec}}$, LLaMA-2.  [^2] Future works applying CEPE to LLaMA-2 models can simply use our pre-trained encoder. We train $\mathcal{M}_{\textrm{enc}}$ for 100K steps with a batch size of $2,048$ and sequence length of $512$ tokens. For more details, refer to §A.1.
@@
We freeze the weights of $\mathcal{M}_{\textrm{dec}}$ and train only the added cross-attention layers as well as fine-tuning $\mathcal{M}_{\textrm{enc}}$ using the cross-entropy loss. We first adopt a warmup training stage designed to teach $\mathcal{M}_{\textrm{dec}}$ to copy from $\mathcal{M}_{\textrm{enc}}$ through cross-attention—for each position $i\leq T$, the objective is to generate $x_{i+1}$, conditioned on $\mathcal{M}_{\textrm{enc}}(x_{1},\ldots,x_{T})$ and $\mathcal{M}_{\textrm{dec}}(x_{1},\ldots,x_{i})$$\mathcal{M}_{\textrm{enc}}$ and $\mathcal{M}_{\textrm{dec}}$ share the same input. The warmup stage uses 131M tokens from the training set of RP. [^3] Preliminary experiments suggest this stage can stabilize training; ablations can be found in §7.1.
@@ 2.4 CEPED for Instruction-Tuned Models
Table 1: Long-context language modeling text perplexity on ArXiv and Book from RedPajama, PG19, ProofPile, and CodeParrot. Throughput compares the speed (number of tokens/second) of each model with that of LLaMA-2. All experiments are conducted on one A100 80GB GPU, except for LLaMA-2-32K and YaRN with 128K tokens, which require model parallelism and are conducted on four A100 GPUs.
@@ 3 Long-context Language Modeling
We focus on 7B parameters models. Our baseline includes LLaMA-2 and its long-sequence fine-tuned versions: LLaMA-2-32K Together (2023a), YaRN-64K, and YaRN-128K Peng et al. (2024). LLaMA-2-32K was trained on RP with upsampled ArXiv and Books domains, and a mixture of other data , and YaRN was trained on PG19. While YaRN-64K and YaRN-128K were fine-tuned for only 1.7B and 3.4B tokens, respectively, fine-tuning all parameters requires much more memory relative to CEPE. We also evaluate on training-free long-context methods: StreamingLLM Xiao et al. (2024b) and RePlug Shi et al. (2023) with LLaMA-2-7B. Although RePlug was originally evaluated in a retrieval-augmented setting, we found that it also works well in long-context modeling, when viewing the long context as retrieved context. See §B for more implementation details.
@@ Results.
We show the results in Table 1. Compared to the two fully fine-tuned models, LLaMA-2-32K and YaRN-64K, CEPE achieves either lower or comparable perplexity across all datasets with lower memory usage and higher throughput despite being trained only 8K sequences. Furthermore, CEPE continues to improve on perplexity while maintaining low memory use at 128K tokens, well beyond its training lengths (8K); on the other hand, LLaMA-2-32K and YaRN-64K cannot generalize beyond its training length and the memory cost increases significantly. At 128K tokens, we outperform or achieve comparable perplexity with all applicable baselines on all domains. Although CEPE observes higher perplexity than YaRN-64K on PG19 and Book at 32K, we note that YaRN-64K was fine-tuned on PG19, and thus has a significant domain advantage.
@@ 4.2 Open-domain Question Answering
LLaMA-2 40 94.3 98.7 74.7 52.3 87.7 54.8 95.1 76.7 62.1 50.4 72.0
Table 3: ICL results averaged across 3 seeds. $k$ is the number of demonstrations. All models use 2 demonstrations in the decoder, and we add $+k^{\prime}$ demonstrations to the encoder for CEPE. denotes the oracle setting with $k=40$ demonstration in the decoder.
@@
Table 4: ZeroScrolls validation results. The total number of tokens includes both the input and generated tokens. For NarrativeQA (NQA) and Qspr (QASPER), we report the F1 scores. For QALT (QuALITY), we report accuracy. For GovReport (GvRp), SummScreenFD (SSFD), and QMSum, we report the ROUGE-L scores. CEPED uses 2K tokens in the decoder, and additional tokens are inputted through the encoder. Only NQA has a substantial number of test examples with more than 32K tokens, so we exclude other dataset on 2K + All setting; see Table 11.
@@ Results.
The results are shown in Figure 3 CEPE consistently outperforms all models across all datasets and $k$. Notably, CEPE outperforms LLaMA-2-32K on NQ and TriviaQA by over 3 and 4 EM points, respectively. Furthermore, CEPE does not degrade in performance as the $k$ increases, while other models often perform worse at larger $k$, as they are sensitive to the large amount of redundant or irrelevant passages Liu et al. (2024).
@@
Table 4 shows that CEPED improves upon LLaMA-2-Chat with 2K tokens across all tasks. The performance of CEPED improves or remains consistent as we scale up the number of tokens in the context window. Notably, on NQA—the only dataset with a significant number of samples longer than 32K tokens (see Table 11)—CEPED improves upon LLaMA-2-Chat by 3 points in F1 scores. Furthermore, CEPED outperforms LLaMA-2-32K Instruct on 4 out of the 6 tasks, despite being trained on unlabeled data. We provide qualitative examples and analysis in §C.3.